OpenStudy (exoexoexo):

if five pencils and three erasers cost 65 cents, and three pencils and two erasers cost 40 cents, what is the cost of one pencil and one eraser?

OpenStudy (anonymous):

ok. wait 2 minutes

OpenStudy (anonymous):

pencil cost 10 cents,eraser5 cents

OpenStudy (exoexoexo):

can you explain to me how you got it?

OpenStudy (anonymous):

yes we do equation 5p+3e=65(multiply on 2 for getting 6e) 3p+2e=40(multimply on 3 to get 6e) we got 10p+6e=130 and 9p+6e=120 6e=130-10p 9p+130-10p=120 -p=-10 p=10

OpenStudy (anonymous):

it is one pencil

OpenStudy (anonymous):

but now we only put in one of equations 50+3e=65 3e=15 e=5 it is eraser